OpenStudy (anonymous):

Okay one more question.. well, more of a review of my work. Anyway... I was told to find (f o g o h)(x). I'll put the rest in the comments.

OpenStudy (anonymous):

I know that (f o g o h)=f(g(h(x))), which seems pretty confusing but I understand it pretty well. Anyway,\[f(x)=3x\]\[g(x)=x-3\]\[h(x)=\sqrt{x}\]so what I did first was find (f o g) which is \[(f o g)=3x-9\] and added h(x) to it to get\[(f o g o h)=3(\sqrt{x})-9\]

OpenStudy (anonymous):

does this need simplifying? if so, how do I simplify it?

OpenStudy (anonymous):

@phi can you check my work, please?

hartnn (hartnn):

your final answer is correct, but we first find g(h(x)) then f( g(h(x)))

OpenStudy (anonymous):

Ohh, okay that makes sense too (:

OpenStudy (anonymous):

So the final answer is correct? It doesn't need to be simplified any further?

OpenStudy (anonymous):

Moreso it CAN'T be simplified any further without the value of the variable?

OpenStudy (phi):

it looks good

hartnn (hartnn):

yeah, can't be simplified..

OpenStudy (anonymous):

Okay thank yous (: